OpenStudy (anonymous):

Natasha Gargarin paid $195 interest on a $2000 purchase. If she made 12 equal monthly payments to pay off the account, what was the APR for this purchase?

OpenStudy (amistre64):

think that last ones a little off.....

OpenStudy (anonymous):

16 1/2% (that is, 16.5%) 17% 17 1/2% (that is, 17.5%) 18%

OpenStudy (anonymous):

thats all i can choose from

OpenStudy (anonymous):

aww u deleted it... i was looking at it lol

OpenStudy (amistre64):

1) simple interest should be: 195 = 2000*12*r r = 195/(24000) r = .8125 % 2) compounded interest would be: 2195 = 2000(1+r/12)^12 2195/2000 = (1+r/12)^12 (2195/2000)^(1/12) = 1 + r/12 (2195/2000)^(1/12) - 1 = r/12 r = 12* [(2195/2000)^(1/12) - 1] maybe lol

OpenStudy (anonymous):

you know i tried to send the ? to KGB and they got it wrong

OpenStudy (amistre64):

that doesnt match an answer either does it... r = 9.33% with that...

OpenStudy (amistre64):

well, getting it wrong is sometimes better than getting it right lol

OpenStudy (anonymous):

see i got 5 questions for this exam and im barley passing so im trying to get all five right lol

OpenStudy (amistre64):

let me do a quick read on APR .... there is something that im forgetting, I just know it ;)

OpenStudy (anonymous):

andthis last question i cant get

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

........ 17.5 is what I can come up with; but how? i dunno

OpenStudy (anonymous):

hey... ima go with it

OpenStudy (amistre64):

when I put in 17.5; I get interest of 194......

OpenStudy (amistre64):

hope its right ;)

OpenStudy (anonymous):

ill let you now in a sec

OpenStudy (anonymous):

We need to use Table 13.3 to determine the APR, and it is displayed based on the finance charge per $100. So we need to calculate how much finance charge Natasha paid per $100. ($195/$2000)*($100) = $9.75 Turning to Table 13.3, for the line with no. payments = 12, $9.75 is closest to the column for 17.5%. So the APR is approximately 17 1/2% (17.5%)

OpenStudy (anonymous):

good job!! thanks alot!!

OpenStudy (amistre64):

9.75 i had .... just couldnt figure what to do with it ;)